OpenStudy (anonymous):

What is the solution to the equation log2 4 – log2 x = 3 ?

OpenStudy (anonymous):

log2 (4/x) = 3 (4/x ) = 2^3 = 8 - can you finish it?

OpenStudy (anonymous):

i used on of laws of logarithms to get the first line: log a - log b = log (a/b)

OpenStudy (anonymous):

no i can't finish it